OpenStudy (anonymous):

Phosphorus-32 has a half-life of 14.0 days. Starting with 8.00 g of ^{32}P, how many grams will remain after 56.0 days?

OpenStudy (anonymous):

0.5 grams. (56 days is 4 half-lifes, yes? therefore, 8/2= 4 /2 = 2 / 2 = 1 /2 = 0.5)